The length of a rectangle is three times its width. If the
length is decreased by 20 ft. and the width increased by 10 ft.,
the area is increased by 200 sq. ft. Fund the dimensions
of the rectangle.

Answer by rothauserc(4718)

length(l) = 3*width(w) and
Area = 3w^2
(3w-20) * (w+10) = 3w^2 + 200
3w^2 +10w -200 = 3w^2 +200
10w = 400
w = 40 feet
l = 120 feet
check answer
100 * 50 = 4800 + 200
5000 = 5000
our answer checks